# Defending Large Language Models Against Attacks With Residual Stream Activation Analysis thanks: Research sponsored by HiddenLayer

### v3

@@ V Results
TABLE VIII: Fine Tuned LLaMA 2 7B Results - Averaged Over Layers
Dataset Accuracy Std Deviation Precision Prompt Class Recall
Broad 0.99941 0.00042 Attack 0.99947 0.99936
Benign 0.99936 0.99947
Domain-Specific 0.98456 0.01170 Attack 0.99323 0.97571
Benign 0.97635 0.99341
Hyper-Specific 0.99976 0.00039 Attack 0.99989 0.99963
Benign 0.99963 0.99989
V-A LLM Activation Analysis - Classification
@@
Tables II, III, and IV show the average classification accuracies across the transformer layers since we found that the statistics were very similar across layers, except for a very slightly lower accuracy in the first five layers and a higher accuracy in the last few layers. Overall it’s clear that this classifier has no difficulty separating the prompt classes. Additionally, the classifier performs equally well on Wizard Vicuna 7B Uncensored as the other safety-tuned models, indicating that the classification success on the other models is not due to any rejection sequences. The classification of the Domain-Specific dataset is slightly worse in all statistics across all models when compared to the other datasets. This is unexpected because the attack prompts in the Domain-Specific dataset are more simplistic and repetitive due to the nature of their creation (details in Appendix A).
@@
We repeat the analysis for our length restricted prompt datasets and compiled the results into tables V, VI, and VII. Even though there was a slight decrease in accuracy in these transformed datasets, the accuracy of the classifier models overall is still high. We include a sample of accuracy scores for each classifier trained on individual layers of LLaMA 2 7B, compared with the original dataset version in Appendix B. This sample shows that every classifier for every layer in the length restricted version of the analysis results in a slightly lower accuracy.
@@
V-B LLM Activation Analysis - Fine Tuned
We repeat our analysis on our fine tuned versions of LLaMA 2 7B. We summarize the results from training a LightGBM classification model in Table VIII where each statistic is the average across all 32 layers of the fine tuned versions of LLaMA 2 7B [5]. The results are very similar to the non-fine tuned version of LLaMA 2 7B and it is not clear at this time that there is a statistically significant difference between the classification results of the models.
VI Conclusion
The classification accuracy of the LightGBM models shows promising results. The accuracy, recall, and precision are consistently high across each dataset type and model. This is significant because the attack behavior varies across dataset type but is still correctly classified. Our prompt length restrictions demonstrate that the classifier is not relying on length as a proxy for prompt class and our random relabeling procedure demonstrates that there is a significant separation between classes in the residual activation representational space. Additionally, since the classification accuracy is high even for classifiers trained on individual layers, it’s possible that any defense based on this technique may only need information from one or two transformer layers to detect attacks.
@@ VI Conclusion
TABLE IX: Classification Accuracy by Layer - LLaMA 2 7B, Broad Attack Dataset
Layer Original Length Restricted Randomly Relabeled
Accuracy Accuracy Accuracy
0 1.00000 0.99886 0.50602
1 1.00000 0.99967 0.50163
2 1.00000 0.99935 0.49365
3 1.00000 0.99919 0.49902
4 1.00000 0.99870 0.49398
5 1.00000 0.99837 0.49333
6 1.00000 0.99951 0.49642
7 1.00000 0.99935 0.50309
8 1.00000 0.99951 0.48779
9 0.99999 0.99967 0.48844
10 1.00000 0.99935 0.49691
11 1.00000 0.99951 0.50081
12 1.00000 0.99967 0.50179
13 1.00000 0.99951 0.49089
14 1.00000 0.99967 0.49642
15 0.99988 0.99951 0.49691
16 0.99994 0.99984 0.49577
17 0.99994 0.99951 0.49919
18 1.00000 0.99984 0.49463
19 1.00000 0.99984 0.49154
20 1.00000 0.99951 0.50033
21 1.00000 0.99951 0.50065
22 1.00000 0.99967 0.48877
23 1.00000 0.99967 0.49186
24 1.00000 0.99967 0.50749
25 1.00000 0.99984 0.50146
26 1.00000 0.99967 0.50293
27 0.99994 0.99984 0.50277
28 1.00000 0.99951 0.50439
29 1.00000 0.99967 0.48812
30 1.00000 0.99984 0.49919
31 1.00000 0.99967 0.49235
@@ Appendix A Domain-Specific Attack Prompt Generation
We utilize Wizard Vicuna Uncensored 7B in order to automate this process since the uncensored nature of this model allows users to make requests of the model without risk of rejection. [18]. We start this process by generating 20 different types of financial harm using the uncensored model and then making minor adjustments to the list in order to differentiate the harms as much as possible. Some of the types of financial harm we use include:
@@ Appendix B Example Classifier Scores Per Layer
Table IX is a sample of the data generated for one model and dataset pairing (LLaMA 2 7B and the Broad attack dataset). For every layer of the model, a separate LightGBM classifier is trained on the activations for that layer. We report the overall accuracy of the original dataset as well as the prompt length range restricted version and the randomly relabeled version. Most of the other model and dataset pairing results are similar in accuracy for all three iterations of every dataset, with the original dataset accuracy scoring close to 1.00000 for every layer, the range restricted version scoring close to 0.999, and the randomly relabeled version never scoring higher than 51% for any layer. The only exception is the model and dataset combination for the Domain-Specific attack dataset, which has slightly lower accuracy for both the original and length range restricted iterations for every layer. Also notably, the later layers of each model generally score slightly higher in accuracy as compared to the first five layers.

### v4

@@ 5 Results
Table 3: Domain-Specific Attack Results - Averaged Over Layers
Model Accuracy±$\pm$±Std. Dev. False Positive Rate False Negative Rate
LLaMA 2 7B 0.992$\pm$0.0056 0.005 0.011
LLaMA 2 13B 0.994$\pm$0.0067 0.002 0.009
TinyLlama 1.1B 0.986$\pm$0.0100 0.009 0.019
Mistral 7B 0.987$\pm$0.0043 0.005 0.020
Vicuna 7B 0.984$\pm$0.0077 0.008 0.027
Wizard Vicuna 7B Uncensored 0.991$\pm$0.0086 0.006 0.016
Table 4: Hyper-Specific Attack Results: HackAPrompt 2023 - Averaged Over Layers
Model Accuracy±$\pm$±Std. Dev. False Positive Rate False Negative Rate
LLaMA 2 7B 0.999$\pm$0.0003 0.001 0.001
LLaMA 2 13B 0.999$\pm$0.0008 0.001 0.001
TinyLlama 1.1B 0.999$\pm$0.0011 0.001 0.001
Mistral 7B 0.999$\pm$0.0002 0.001 0.001
Vicuna 7B 0.999$\pm$0.0005 0.001 0.001
Wizard Vicuna 7B Uncensored 0.999$\pm$0.0008 0.001 0.001
5.1 LLM Activation Analysis - Classification
Results for Non-Length Restricted Prompts
@@ 5.1 LLM Activation Analysis - Classification
Tables 2, 3, and 4 show the average classification accuracies across the transformer layers since we found that the statistics were very similar across layers, except for a very slightly lower accuracy in the first five layers and a higher accuracy in the last few layers. Overall it’s clear that this classifier has no difficulty separating the prompt classes for these datasets. These results make sense, given that LLMs can act as high-performance feature extractors [6].
Additionally, the classifier performs equally well on Wizard Vicuna 7B Uncensored as the other safety-tuned models, indicating that the classification success on the other models is not due to any rejection sequences. The classification of the Domain-Specific dataset is slightly worse in all statistics across all models when compared to the other datasets. This is unexpected because the attack prompts in the Domain-Specific dataset are more simplistic and repetitive due to the nature of their creation (details in Appendix A). However, a limitation not reflected in the tables is that in the experiments on LLaMA 2 7B, we found that a classifier trained on the pre-processing only activations (as described in 3) performs just as well as all other classifiers trained on transformer activations. An example of this result is shown in Appendix B as “layer 0”.
Table 5: Broad Attack Results (Length Range Restricted) - Averaged Over Layers
Model Accuracy±$\pm$±Std. Dev. False Positive Rate False Negative Rate
LLaMA 2 7B 0.999$\pm$0.0003 0.001 0.001
LLaMA 2 13B 0.999$\pm$0.0006 0.001 0.001
TinyLlama 1.1B 0.999$\pm$0.0007 0.001 0.001
Mistral 7B 0.999$\pm$0.0003 0.001 0.001
Vicuna 7B 0.999$\pm$0.0004 0.001 0.001
Wizard Vicuna 7B Uncensored 0.999$\pm$0.0004 0.001 0.001
Table 6: Domain-Specific Attack Results (Length Range Restricted) - Averaged Over Layers
Model Accuracy±$\pm$±Std. Dev. False Positive Rate False Negative Rate
LLaMA 2 7B 0.991$\pm$0.0060 0.006 0.013
LLaMA 2 13B 0.993$\pm$0.0066 0.004 0.011
TinyLlama 1.1B 0.990$\pm$0.0085 0.009 0.011
Mistral 7B 0.990$\pm$0.0034 0.007 0.013
Vicuna 7B 0.992$\pm$0.0055 0.005 0.011
Wizard Vicuna 7B Uncensored 0.990$\pm$0.0074 0.005 0.014
Table 7: Hyper-Specific Attack Results: HackAPrompt 2023 (Length Range Restricted) - Averaged Over Layers
Model Accuracy±$\pm$±Std. Dev. False Positive Rate False Negative Rate
LLaMA 2 7B 0.999$\pm$0.0012 0.001 0.002
LLaMA 2 13B 0.999$\pm$0.0027 0.001 0.002
TinyLlama 1.1B 0.999$\pm$0.0016 0.002 0.001
Mistral 7B 0.999$\pm$0.0006 0.001 0.001
Vicuna 7B 0.999$\pm$0.0011 0.001 0.001
Wizard Vicuna 7B Uncensored 0.999$\pm$0.0015 0.001 0.001
As seen from the results in Table 8, the detection methodology did not perform as well on the WildJailbreak dataset. This is expected, as the difference in prompt classes is not as well defined as in the other datasets. Table 12 in Appendix B provides a breakdown of classifier performance by layer of LLaMA 2 7B on this dataset. Performance improves up until the layers midway through the model with no significant improvement in the following layers. Notably, the classifier trained on the preprocessing-only layer, “layer 0,” performs slightly better than the classifiers trained on the first few layers of the model.
As seen from the results of the holdout test in Table 9, the classifier is unable to correctly classify the unseen attack type, regardless of LLM used. The high false negative rate indicates the tendency of the classifier to mistake the unseen attack type as a benign prompt.
Results for Length Restricted Prompts
@@
We repeat the analysis for our length restricted prompt datasets and compiled the results into tables 5, 6, and 7. Even though there was a slight decrease in accuracy in these transformed datasets, the accuracy of the classifier models overall is still high. This discrepancy in performance is most likely due to the reduced dataset size for the length restricted versions. We hypothesize that there would be no discrepancy in performance if the non-restricted and restricted datasets had the same size. We include a sample of accuracy scores for each classifier trained on individual layers of LLaMA 2 7B, compared with the original dataset version in Appendix B. This sample shows that every classifier for every layer in the length restricted version of the analysis results in a slightly lower accuracy.
Results for Random Relabeling
@@
Table 8: Hyper-Specific Attack Results: WildJailBreak - Averaged Over Layers
Model Accuracy±$\pm$±Std. Dev. False Positive Rate False Negative Rate
LLaMA 2 7B 0.825$\pm$0.059 0.268 0.083
LLaMA 2 13B 0.854$\pm$0.067 0.216 0.075
TinyLlama 1.1B 0.740$\pm$0.024 0.411 0.108
Mistral 7B 0.820$\pm$0.046 0.272 0.087
Vicuna 7B 0.830$\pm$0.059 0.263 0.078
Wizard Vicuna 7B Uncensored 0.830$\pm$0.056 0.261 0.079
Table 9: Broad Attack Results Validated on Holdout Attack Style - Averaged Over Layers
Model Accuracy±$\pm$±Std. Dev. False Positive Rate False Negative Rate
LLaMA 2 7B 0.581$\pm$0.051 0.012 0.827
LLaMA 2 13B 0.594$\pm$0.053 0.010 0.803
TinyLlama 1.1B 0.633$\pm$0.084 0.012 0.722
Mistral 7B 0.573$\pm$0.064 0.011 0.843
Vicuna 7B 0.549$\pm$0.012 0.078 0.886
Wizard Vicuna 7B Uncensored 0.550$\pm$0.046 0.012 0.887
5.2 LLM Activation Analysis - Fine Tuned
We repeat our analysis on our fine tuned versions of LLaMA 2 7B. We summarize the results from training a LightGBM classification model in Table 10 where each statistic is the average across all 32 layers of the fine tuned versions of LLaMA 2 7B [5]. The results are very similar to the non-fine tuned version of LLaMA 2 7B and it is not clear at this time that there is a statistically significant difference between the classification results of the models.
Table 10: Fine Tuned LLaMA 2 7B Results - Averaged Over Layers
Dataset Accuracy±$\pm$±Std. Dev. False Positive Rate False Negative Rate
Broad 0.999$\pm$0.0004 0.001 0.001
Domain-Specific 0.985$\pm$0.0117 0.006 0.013
Hyper-Specific 0.999$\pm$0.0004 0.001 0.001
6 Conclusion
The classification accuracy of the LightGBM models on activations shows promising results. For most datasets, the classifier consistently had high accuracy, low false positive, and low false negative rates across all layers of the LLMs. This is significant because high performance is consistent across attack types, regardless of attack type. Our prompt length restrictions demonstrate that the classifier is not relying on length as a proxy for class and our random relabeling procedure demonstrates that there is a significant separation between classes in the residual activation representational space. This confirms that these decoder-only LLMs can be used as high-performance feature extractors for text data for classification purposes. However, it is unclear that this methodology for feature extraction is necessary in every case. Our limited testing shows that classifiers trained on pre-processing activations perform just as well, except for when there is significant overlap in prompt style and topic between the classes. This indicates that our methodology is most useful when the prompt classes are not easily distinguishable. Additionally, it is important to note that this methodology does not generalize to unseen attacks; that is, attack styles not provided in the training set for the classifier.
@@ 6 Conclusion
Acknowledgements.
The authors would like to acknowledge the valuable feedback Professor Stefan Lee and Professor Sanghyun Hong provided during the development of this research. Research sponsored by HiddenLayer
@@ Appendix A Domain-Specific Attack Prompt Generation
We utilize Wizard Vicuna Uncensored 7B in order to automate this process since the uncensored nature of this model allows users to make requests of the model without risk of rejection [22]. We start this process by generating 20 different types of financial harm using the uncensored model and then making minor adjustments to the list in order to differentiate the harms as much as possible. Some of the types of financial harm we use include:
@@ Appendix B Example Classifier Scores Per Layer
Table 11: Classification Accuracy by Layer - LLaMA 2 7B, Broad Attack Dataset
Layer Original Length Restricted Randomly Relabeled
Accuracy Accuracy Accuracy
0 1.000 0.999 0.555
1 1.000 0.999 0.506
2 1.000 0.999 0.502
3 1.000 0.999 0.494
4 1.000 0.999 0.499
5 1.000 0.999 0.494
6 1.000 0.999 0.493
7 1.000 0.999 0.496
8 1.000 0.999 0.503
9 1.000 0.999 0.488
10 0.999 0.999 0.488
11 1.000 0.999 0.497
12 1.000 0.999 0.501
13 1.000 0.999 0.502
14 1.000 0.999 0.491
15 1.000 0.999 0.496
16 0.999 0.999 0.497
17 0.999 0.999 0.496
18 0.999 0.999 0.499
19 1.000 0.999 0.495
20 1.000 0.999 0.492
21 1.000 0.999 0.500
22 1.000 0.999 0.501
23 1.000 0.999 0.489
24 1.000 0.999 0.492
25 1.000 0.999 0.507
26 1.000 0.999 0.501
27 1.000 0.999 0.503
28 0.999 0.999 0.503
29 1.000 0.999 0.504
30 1.000 0.999 0.488
31 1.000 0.999 0.499
32 1.000 0.999 0.492
Table 11 is a sample of the data generated for one model and dataset pairing (LLaMA 2 7B and the Broad attack dataset). For every layer of the model, a separate LightGBM classifier is trained on the activations for that layer. We report the overall accuracy of the original dataset as well as the prompt length range restricted version and the randomly relabeled version. Most of the other model and dataset pairing results are similar in accuracy for all three iterations of every dataset, with the original dataset accuracy scoring close to 1.00000 for every layer, the range restricted version scoring close to 0.999, and the randomly relabeled version never scoring higher than 51% for any layer. The only exception is the model and dataset combination for the Domain-Specific attack dataset, which has slightly lower accuracy for both the original and length range restricted iterations for every layer. Also notably, the later layers of each model generally score slightly higher in accuracy as compared to the first five layers. Additionally, the zeroth layer classifier, which corresponds to the activations after pre-processing the prompt, generally performs just as well as the classifiers for the other layers. The exception to this pattern is shown in Table 12, which demonstrates how classifiers for the mid-layer activations outperform the pre-processing layer classifier.
Table 12: Analysis by Layer, LLaMA 2 7B, WildJailBreak Dataset
Layer Accuracy±$\pm$±Std. Dev. False Positive Rate False Negative Rate
0 0.732 0.372 0.164
1 0.690 0.434 0.186
2 0.712 0.446 0.130
3 0.719 0.450 0.112
4 0.714 0.440 0.132
5 0.755 0.380 0.110
6 0.747 0.384 0.122
7 0.762 0.364 0.112
8 0.794 0.318 0.094
9 0.813 0.298 0.076
10 0.830 0.250 0.090
11 0.845 0.232 0.078
12 0.832 0.270 0.066
13 0.868 0.208 0.056
14 0.848 0.238 0.066
15 0.872 0.194 0.062
16 0.871 0.184 0.074
17 0.846 0.214 0.094
18 0.883 0.178 0.056
19 0.867 0.216 0.050
20 0.874 0.198 0.054
21 0.872 0.196 0.060
22 0.876 0.198 0.050
23 0.866 0.208 0.060
24 0.868 0.198 0.066
25 0.859 0.212 0.070
26 0.848 0.236 0.068
27 0.874 0.198 0.054
28 0.853 0.238 0.056
29 0.862 0.224 0.052
30 0.860 0.212 0.068
31 0.873 0.196 0.058
32 0.833 0.256 0.078